The sum of the base and the height of a triangle is 20cm. Find the dimension for which the area is a maximum
Answer by checkley71(8403)

AREA=(B*H)/2
INORDER FOR THE AREA TO BE MAXED THE B & H SHOULD BE THE SAME.
SINCE THE SUM IS 20 THEN 20/2=10
AREA=(10*10)2
AREA=100/2
AREA=50CM.